# Statistical description and dimension reduction of categorical trajectories with multivariate functional principal components

### v3

@@ 2 Notations and mathematical framework
setting to one the trace of the covariance operator $(w_{j})^{-1}\Gamma_{jj}$ of the binary (normalized) trajectory $X_{j}$.
@@
Previous framework can be extended easily to experiments in which more than two states can be present simultaneously without increasing the cardinality of $\mathcal{S}$ (see the TCATA experiment described in Section 4.2). If we replace $\mathbf{H}_{1}$ by the hypothesis that the trajectories $X_{1},\ldots,X_{q}$ are continuous in probability, Proposition 2.1 remains true.
3 Sampling and estimators
Suppose now we have a sample of $n$ categorical trajectories $Y_{1},\ldots,Y_{n}$ observed over $[0,T]$ and taking values in $\mathcal{S}=\{S_{1},\ldots,S_{q}\}$. We define $\mathbf{X}_{i}=(X_{i1},\ldots,X_{iq})$ the vector of the $q$ corresponding binary trajectories, with $X_{ij}(t)=\mathbf{1}_{\{Y_{i}(t)=S_{j}\}}$. These trajectories are piecewise constant, with a finite number of jumps. It is a rare case in functional data analysis in which the trajectories can be observed exhaustively, that is to say for all time point $t\in[0,T]$.
We consider, for all $s$ and $t$ in $[0,T]$, the empirical probabilities of occurrence $\widehat{p}_{j}(t)=n^{-1}\sum_{i=1}^{n}X_{ij}(t)$ and $\widehat{\boldsymbol{p}}(t)=(\widehat{p}_{1}(t),\ldots,\widehat{p}_{q}(t))$. We also consider the joint empirical probabilities $\widehat{p}_{j\ell}(s,t)=n^{-1}\sum_{i=1}^{n}X_{ij}(s)X_{i\ell}(t)$, and define the estimators of the covariance functions
$\widehat{\gamma}_{j}(t,t)=\widehat{p}_{j}(t)(1-\widehat{p}_{j}(t))$
and
$\widehat{\gamma}_{j\ell}(s,t)=\widehat{p}_{j\ell}(s,t)-\widehat{p}_{j}(s)\widehat{p}_{\ell}(t).$
We deduce from (2), estimators $\widehat{\Gamma}_{j\ell}$ of the cross-covariance operators $\Gamma_{j\ell}$ and from (5), an estimator $\widehat{\boldsymbol{\Gamma}}$ of $\boldsymbol{\Gamma}$. Note that in a more formal way, we can express
$\widehat{\boldsymbol{p}}$ $=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{X}_{i}$
$\widehat{\boldsymbol{\Gamma}}$ $=\sum_{i=1}^{n}\boldsymbol{X}_{i}\otimes\boldsymbol{X}_{i}-\widehat{\boldsymbol{p}}\otimes\widehat{\boldsymbol{p}}$
where the tensor product is defined as follows, $\left(\boldsymbol{u}\otimes\boldsymbol{v}\right)\boldsymbol{\phi}=\langle\boldsymbol{u},\boldsymbol{\phi}\rangle_{\mathbb{H}}\boldsymbol{v}$ for all $\boldsymbol{u},\boldsymbol{v},\boldsymbol{\phi}$ belonging to $\mathbb{H}$. We can state the following consistency and asymptotic normality results which follow immediately from assumption $\mathbf{H}_{1}$. We denote by $\|\boldsymbol{\Gamma}\|_{op}$ the spectral norm of linear operator $\boldsymbol{\Gamma}$ induced by the norm in $\mathbb{H}$.
@@ 3 Sampling and estimators
Suppose that hypothesis $\mathbf{H}_{1}$ is fulfilled, as $n$ tends to infinity,
  • $\left\|\widehat{\mathbf{p}}-\mathbf{p}\right\|_{\mathbb{H}}\to 0$ and $\left\|\widehat{\boldsymbol{\Gamma}}-\boldsymbol{\Gamma}\right\|_{op}\to 0$ almost surely,
  • $\sqrt{n}\left(\widehat{\mathbf{p}}-\mathbf{p}\right)$ and $\sqrt{n}\left(\widehat{\boldsymbol{\Gamma}}-\boldsymbol{\Gamma}\right)$ are asymptotically Gaussian.
@@
First note that $\mathbb{H}$ equipped with the inner product $\langle.,.\rangle_{\mathbb{H}}$ is a separable Hilbert space. We clearly have
$\|\mathbf{X}\|_{\mathbb{H}}^{2}$ $=\sum_{j=1}^{q}w_{j}\|X_{j}\|^{2}$
$\leq T\max_{j}w_{j},$
@@
We can deduce from Proposition 3.1 the consistency and asymptotic normality of the estimators of the eigenvalues $\widehat{\lambda}_{r}$, $r=1,\ldots,k$ (see Dauxois et al., (1982) or Theorem 9.1.3 in Hsing and Eubank, (2015)). Under the additional assumption that the eigenvalues are distinct, we can also get the asymptotic normality of the estimated eigenfunctions $\widehat{\boldsymbol{\phi}}_{r}(t)$, $r=1,\ldots,k$.
4 Two illustrations in sensory analysis
The dataset used for illustration deals with sensory experiments. It is well documented in Béno et al., (2023) and can be obtained from a public source. In each illustration fifty panelists took part in a tasting experiment and were asked to click on the sensation they perceived in real time from a list of descriptors. Computations were performed with the library MFPCA (see Happ, (2022)) in the R language R Core Team, (2024). All codes are available on Github, https://github.com/Chemosens/ExternalCode/tree/main/MFPCAWithCategoricalTrajectories.

### v4

@@ 2.1 The continuous time stochastic process point of view
setting to one the trace of the covariance operator $(w_{j})^{-1}\Gamma_{jj}$ of the binary (normalized) trajectory $X_{j}$ (see the supplementary material for an illustration on the sensory data).
2.2 The functional point of view and multivariate PCA
We now adopt another point of view and introduce the space $D([0,T],\mathbb{R}^{q})$, a multidimensional version of the space $D([0,T])$, the set of functions which have only jump discontinuities and are right continuous (see Chapter 14 in Breiman,, 1968 or Chapter 3 in Billingsley,, 1999).
Definition 1.
Let $D([0,T],\mathbb{R}^{q})$ be the set of functions $\mathbf{x}:[0,T]\mapsto\mathbb{R}^{q}$ that are right-continuous and have left-hand limits,
  • for $0\leq t<T$, $\mathbf{x}(t+)=\lim_{s\searrow t}\mathbf{x}(s)$ exists and $\mathbf{x}(t+)=\mathbf{x}(t)$,
  • for $0<t\leq T$, $\mathbf{x}(t-)=\lim_{s\nearrow t}\mathbf{x}(s)$ exists.
As stated in Breiman, (1968), an element $\mathbf{x}$ of $D([0,T],\mathbb{R}^{q})$ has at most countably many jumps, and for any $\delta>0$ and the set $\{t\in[0,T];\|\mathbf{x}(t)-\mathbf{x}(t-)\|\geq\delta\}$ is finite, where $\|.\|$ is a norm on $\mathbb{R}^{q}$.
The following proposition allows to get a joint measurability result under the weak hypothesis that the trajectories belong to $D([0,T],\mathbb{R}^{q})$. Note that it is a general statement, that does not assume that the trajectories are drawn from TDS or TCATA experiments. It can be seen as a generalization of Theorem 7.4.2 in Hsing and Eubank, (2015) which replaces the hypothesis that the trajectories $X(.,\omega)$ are continuous for all $\omega\in\Omega$ by the slightly weaker assumption that they belong to $D([0,T],\mathbb{R}^{q})$.
Proposition 2.2.
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and let $\mathbf{Z}:[0,T]\times\Omega\rightarrow\mathbb{R}^{q}$ be a stochastic process such that
  • for every $t\in[0,T]$, the random variable $\omega\mapsto\mathbf{Z}(t,\omega)$ is $\mathcal{F}$-measurable,
  • for every $\omega\in\Omega$, the trajectory $\mathbf{Z}(.,\omega):t\in[0,T]\mapsto\mathbf{Z}(t,\omega)$ belongs to $D([0,T],\mathbb{R}^{q})$.
Then $\mathbf{Z}$ is jointly measurable, that is,
$\mathbf{Z}:([0,T]\times\Omega,\ \mathcal{B}([0,T])\otimes\mathcal{F})\rightarrow(\mathbb{R}^{q},\ \mathcal{B}(\mathbb{R}^{q}))$
is measurable.
Proof.
The proof follows the same steps as the proof of Theorem 7.4.2 in Hsing and Eubank, (2015), which is stated for continuous trajectories. Let $(q_{k})_{k\geq 1}$ be an increasing enumeration of $\mathbb{Q}\cap[0,T]$ with $q_{1}=0$ and $q_{k}\uparrow T$. For each $n\geq 1$, choose a rational partition
$0=q_{1}<q_{2}<\cdots<q_{N_{n}}=1\quad\text{such that}\quad\max_{1\leq k\leq N_{n}-1}(q_{k+1}-q_{k})\leq 2^{-n},$
and consider the approximation process
$\mathbf{Z}_{(n)}(t,\omega)=\sum_{k=1}^{N_{n}-1}\mathbf{Z}(q_{k},\omega)\,\mathbf{1}_{[q_{k},q_{k+1})}(t)+\mathbf{Z}(T,\omega)\,\mathbf{1}_{\{T\}}(t).$
For each $k$, the mapping $(t,\omega)\mapsto\mathbf{Z}(q_{k},\omega)\,\mathbf{1}_{[q_{k},q_{k+1})}(t)$ is measurable with respect to $\mathcal{B}([0,T])\otimes\mathcal{F}$, since $t\mapsto\mathbf{1}_{[q_{k},q_{k+1})}(t)$ is Borel measurable and $\omega\mapsto\mathbf{Z}(q_{k},\omega)$ is $\mathcal{F}$-measurable. Hence $\mathbf{Z}_{(n)}$ is jointly measurable.
Fix $(t,\omega)\in[0,T]\times\Omega$. There exists an index $k_{n}$ such that $q_{k_{n}}\leq t<q_{k_{n}+1}$. By right-continuity of the trajectories belonging to $D([0,T],\mathbb{R}^{q})$, $t\mapsto\mathbf{Z}(t,\omega)$, we obtain, as $n$ tends to infinity, $\mathbf{Z}(q_{k_{n}},\omega)\rightarrow\mathbf{Z}(t,\omega)$. Therefore,
$\mathbf{Z}_{(n)}(t,\omega)\rightarrow\mathbf{Z}(t,\omega).$
Since the pointwise limit of a sequence of measurable functions with values in $\mathbb{R}^{q}$ is measurable, it follows that $\mathbf{Z}$ is jointly measurable.
Under hypothesis $\mathbf{H}_{1}$ the components $p_{j}$ are continuous functions and $\mathbf{p}=(p_{1},\ldots,p_{q})$ belongs to $\mathbb{H}$. Next proposition states that the continuous time stochastic process $\mathbf{X}(t)$, that is continuous in probability and has its trajectories in $D([0,1],\mathbb{R}^{q})$, can also be seen as a random element $\mathbf{X}$ taking values in $\mathbb{H}$. Furthermore, it is possible to identify, its expectation in $\mathbb{H}$ with $\mathbf{p}$ and its covariance operator with the integral operator $\boldsymbol{\Gamma}$ defined in (4). We denote by $\left(\boldsymbol{u}\otimes\boldsymbol{v}\right)\boldsymbol{\phi}=\langle\boldsymbol{u},\boldsymbol{\phi}\rangle_{\mathbb{H}}\boldsymbol{v}$ the tensor product of two elements of $\boldsymbol{u}$ and $\boldsymbol{v}$ of $\mathbb{H}$.
Proposition 2.3.
Under hypothesis $\mathbf{H}_{1}$ and if, for all $\omega\in\Omega$, $\mathbf{X}(.,\omega)$ belongs to $D([0,T],\mathbb{R}^{q})$, we have
  • $\mathbb{E}(\mathbf{X})=\mathbf{p}$ in $\mathbb{H}$,
  • $\mathbb{E}\left[\left(\mathbf{X}-\mathbf{p}\right)\otimes\left(\mathbf{X}-\mathbf{p}\right)\right]=\boldsymbol{\Gamma}$.
Proof.
First note that $D([0,T],\mathbb{R}^{q})\subset\mathbb{H}$. By Proposition 2.2, $\mathbf{X}(t,\omega)$ is jointly measurable, and thus $\mathbf{X}$ is a random element in $\mathbb{H}$. The result is then a consequence of hypothesis $\mathbf{H}_{1}$ and Theorem 7.4.3 in Hsing and Eubank, (2015).
We get with Proposition 2.3 that the operator $\boldsymbol{\Gamma}$, defined in (4), is the integral operator with multidimensional kernel function $\boldsymbol{\gamma}(s,t)=(\gamma_{j\ell}(s,t))_{j,\ell=1,\ldots,q}$. For $r\geq 1$, $\boldsymbol{\phi}_{r}$ is an eigenvector associated to the non negative eigenvalue $\lambda_{r}$, $\boldsymbol{\Gamma}\boldsymbol{\phi}_{r}=\lambda_{r}\boldsymbol{\phi}_{r}$. Furthermore, the total variance satisfies
$\text{tr}(\boldsymbol{\Gamma})=\sum_{r\geq 1}\lambda_{r}=\sum_{j=1}^{q}\int_{0}^{T}\gamma_{j,j}(t,t)dt=\mathbb{E}\left[\|\mathbf{X}-\mathbf{p}\|_{\mathbb{H}}^{2}\right]<\infty.$
$\widetilde{\mathbf{X}}_{k}(t)$ $=\mathbf{p}+\sum_{r=1}^{k}\langle\mathbf{X}-\mathbf{p},\boldsymbol{\phi}_{r}\rangle_{\mathbb{H}}\boldsymbol{\phi}_{r}(t),\quad\forall t\in[0,T].$ (11)
3 Sampling and estimators
Suppose now we have a sample of $n$ independent and identically distributed categorical stochastic processes $Y_{1},\ldots,Y_{n}$ observed over $[0,T]$ and taking values in $\mathcal{S}=\{S_{1},\ldots,S_{q}\}$. For $i=1,\ldots,n$, we denote by $\mathbf{X}_{i}=(X_{i1},\ldots,X_{iq})$ the vector of the $q$ corresponding binary trajectories, with $X_{ij}(t)=\mathbf{1}_{\{Y_{i}(t)\ni S_{j}\}}$. These trajectories are piecewise constant, with a finite number of jumps, and belongs to $D([0,T],\mathbb{R}^{q})$. It is a rare case in functional data analysis in which the trajectories can be observed exhaustively, that is to say for all time point $t\in[0,T]$. As in Section 2, we can adopt two different perspectives to study properties of estimators based on $\mathbf{X_{1}},\ldots,\mathbf{X}_{n}$, that of continuous time stochastic processes and that of functional data analysis.
3.1 Stochastic process point of view
We define, for all $t$ in $[0,T]$, the empirical probabilities of occurrence
$\widehat{p}_{j}(t)$ $=\frac{1}{n}\sum_{i=1}^{n}X_{ij}(t),$ (12)
and $\widehat{\boldsymbol{p}}(t)=(\widehat{p}_{1}(t),\ldots,\widehat{p}_{q}(t))$. We also define, for all $s$ and $t$ in $[0,T]$, the joint empirical probabilities
$\widehat{p}_{j\ell}(s,t)=n^{-1}\sum_{i=1}^{n}X_{ij}(s)X_{i\ell}(t).$
The estimators of the covariance functions are $\widehat{\gamma}_{j}(t,t)=\widehat{p}_{j}(t)(1-\widehat{p}_{j}(t))$ and
$\widehat{\gamma}_{j\ell}(s,t)$ $=\widehat{p}_{j\ell}(s,t)-\widehat{p}_{j}(s)\widehat{p}_{\ell}(t).$ (13)
Note that even $p_{j}$ and $\gamma_{j\ell}$ are continuous functions under hypothesis $\mathbf{H}_{1}$, the estimators $\widehat{p}_{j}(t)$ and $\widehat{\gamma}_{j\ell}(s,t)$ have discontinuity points. As stated in the following property, the estimators $\widehat{p}_{j}(t)$ and $\widehat{\gamma}_{j\ell}(s,t)$ converge pointwise almost surely to $p_{j}(t)$ and $\gamma_{j\ell}(s,t)$
Proposition 3.1.
Suppose that $H_{1}$ is true. As $n$ tends to infinity, we have for all $j\in\{1,\ldots,q\}$ and all $t\in[0,T]$, $\widehat{p}_{j}(t)\to p_{j}(t)$ almost surely.
We also have, for all $\ell\in\{1,\ldots,q\}$ and all $s\in[0,T]$, $\widehat{\gamma}_{j\ell}(s,t)\to\gamma_{j\ell}(s,t)$ almost surely.
The proof of Proposition 3.1 follows directly from the continuous mapping theorem and the strong law of large numbers for independent and identically distributed bounded random variables and is therefore omitted.
We can also get uniform convergence of the estimators of the mean functions $p_{j}(.)$ using similar arguments as Yao et al., (2005), Li and Hsing, (2011) or Hsing and Eubank, (2015), under additional smoothness assumptions on the mean trajectory. Furthermore, it follows that observing the trajectories for all $t\in[0,T]$ is not strictly required. Instead, a sufficiently dense grid of $r_{n}$ discretization points, combined with a nonparametric smoothing procedure allow to get parametric rates of converge, up to a logarithmic factor, according to uniform convergence, for the mean under smoothness regularity conditions.
We denote by $T$ a random variable with values in $[0,T]$ and strictly positive density function $f_{T}(.)$ on its support $[0,T]$. We first draw randomly $r_{n}$ observational points, $T_{1},\ldots,T_{r_{n}}$ in $[0,T]$, by considering $r_{n}$ independent copies of $T$. We denote by $X_{ij,k}$ the value of $X_{ij}(T_{k})$ for $j=1,\ldots,q$, $i=1,\ldots,n$ and $k=1,\ldots,r_{n}$. We observe, for $j=1,\ldots,q$ and $i=1,\ldots,n$,
$X_{ij,k}=p_{j}(T_{k})+\left(X_{ij}(T_{k})-p_{j}(T_{k})\right),\quad k=1,\ldots,r_{n}.$ (14)
Estimation of $p_{j}(t)$ can be performed via local linear smoothers, as in Li and Hsing, (2011). Let $K(.)$ be a symmetric probability density function on [1,1][-1,1] and $K_{h}(t)=\frac{1}{h}K\left(\frac{t}{h}\right)$, where $h>0$ is the bandwidth. An estimator of the mean function $p_{j}(t)$ is given by $\widehat{p}_{j}(t)=\widehat{a}_{j,0}$, where
$(\widehat{a}_{j,0},\widehat{a}_{j,1})$ $=\arg\min_{a_{0},a_{1}}\frac{1}{nr_{n}}\sum_{k=1}^{r_{n}}\sum_{i=1}^{n}\left[X_{ij,k}-a_{0}-a_{1}(T_{k}-t)\right]^{2}K_{h_{p}}\left(T_{k}-t\right).$ (15)
Note that in our specific context, and in contrast to Yao et al., (2005), the observed values in (14) are not corrupted by a random noise $\epsilon$. The following property is a direct consequence of Theorem 3.1 and Corollary 3.2 in Li and Hsing, (2011). We omit the proof, noting that the moment conditions (C5) are satisfied since $|X_{ij,k}|\leq 1$.
Proposition 3.2.
Suppose that $0<m_{T}\leq f_{T}(t)\leq M_{T}<\infty$ for all $t\in[0,T]$ and that $f_{T}$ is differentiable with a bounded derivative. Suppose that the kernel function $K(.)$ is a symmetric probability function on [1,1][-1,1], and is of bounded variation on [1,1][-1,1]. We also assume that $p_{j}(.)$ is twice differentiable and the second derivative is bounded on $[0,T]$.
Then if $(h_{p}r_{n})^{-1}=O(1)$ and $h_{p}=O\left((\log n/n)^{1/4}\right)$ as $n$ tends to infinity, we have
$\sup_{t\in[0,T]}\left|\widehat{p}_{j}(t)-p_{j}(t)\right|=O(\sqrt{\log n/n})\quad\text{ almost surely.}$
We deduce that if the number of discretization points $r_{n}$ is sufficiently large, that is to say $n^{1/4}=O(r_{n})$, parametric rates for uniform convergence can be obtained for the estimators of $p_{1},\ldots,p_{q}$, up to a logarithmic term.
@@ 3.1 Stochastic process point of view
Note that similar uniform consistency results as those stated in Proposition 3.2, that do not require smoothness conditions on the individual trajectories, could also be obtained for estimators based on penalized regression splines. See Xiao, (2020) for details on the construction of the estimators of the mean function, as well as asymptotic uniform convergence properties.
Remark 4.
The consistent estimation of the covariance function is not directly possible by such smoothing approaches since the required differentiability assumptions are not fulfilled. Stochastic processes with values in $\{0,1\}$ are not mean square differentiable (see the example given in Appendix C in the Supplementary Material). This lack of regularity translates into non differentiability of the covariance function near the diagonal $(t,t)$, as discussed in Chapter 11, Loève, (1978).
3.2 Functional data point of view
We deduce from (12), (13) and Proposition 2.3, the estimators $\widehat{\Gamma}_{j\ell}$ of the cross-covariance operators $\Gamma_{j\ell}$ and an estimator $\widehat{\boldsymbol{\Gamma}}$ of $\boldsymbol{\Gamma}$. Note that in a more formal way, we can express
$\widehat{\boldsymbol{p}}$ $=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{X}_{i},$
$\widehat{\boldsymbol{\Gamma}}$ $=\sum_{i=1}^{n}\boldsymbol{X}_{i}\otimes\boldsymbol{X}_{i}-\widehat{\boldsymbol{p}}\otimes\widehat{\boldsymbol{p}}$.
We can state the following consistency and asymptotic normality results which follow immediately from assumption $\mathbf{H}_{1}$ and Proposition 2.3. We denote by $\text{HS}(\mathbb{H})$ the vector space of Hilbert-Schmidt operators on $\mathbb{H}$, by $\langle.,.\rangle_{\text{HS}}$ the inner product in $\text{HS}(\mathbb{H})$ and $\Gamma\otimes_{\text{HS}}\Delta=\langle\Gamma,.\rangle_{\text{HS}}\Delta$.
@@ 3.2 Functional data point of view
Suppose that hypothesis $\mathbf{H}_{1}$ is fulfilled and that for almost all $\omega\in\Omega$, $\mathbf{X}(.,\omega)$ belongs to $D([0,T],\mathbb{R}^{q})$, then as $n$ tends to infinity,
  • $\left\|\widehat{\mathbf{p}}-\mathbf{p}\right\|_{\mathbb{H}}\to 0$ almost surely, and $\sqrt{n}\left(\widehat{\mathbf{p}}-\mathbf{p}\right)\overset{d}{\longrightarrow}Z$ in $\mathbb{H}$, where $Z$ is a Gaussian random element with mean zero and covariance operator $\boldsymbol{\Gamma}$.
  • $\left\|\widehat{\boldsymbol{\Gamma}}-\boldsymbol{\Gamma}\right\|_{\text{HS}}\to 0$ almost surely, and $\sqrt{n}\left(\widehat{\boldsymbol{\Gamma}}-\boldsymbol{\Gamma}\right)\overset{d}{\longrightarrow}\mathcal{Z}$ in $\text{HS}(\mathbb{H})$, where $\mathcal{Z}$ is a Gaussian random element with mean zero and covariance operator
    $\mathbb{E}\left[\left((\mathbf{X}-\mathbf{p})\otimes(\mathbf{X}-\mathbf{p})-\boldsymbol{\Gamma}\right)\otimes_{\text{HS}}\left((\mathbf{X}-\mathbf{p})\otimes(\mathbf{X}-\mathbf{p})-\boldsymbol{\Gamma}\right)\right].$
@@
First note that $\mathbb{H}$ equipped with the inner product $\langle.,.\rangle_{\mathbb{H}}$ is a separable Hilbert space, thanks to Proposition 2.3, $\mathbf{X}$ is well defined as a random elements in $\mathbb{H}$. We clearly have, for each $\omega\in\Omega$,
$\|\mathbf{X}(.,\omega)\|_{\mathbb{H}}^{2}$ $=\sum_{j=1}^{q}\|X_{j}(.,\omega)\|^{2}$
$\leq qT,$
@@
We can deduce from Proposition 3.3 the consistency and the asymptotic normality of the estimators of the eigenvalues $\widehat{\lambda}_{r}$, $r=1,\ldots,k$ (see Dauxois et al.,, 1982 or Theorem 9.1.3 in Hsing and Eubank,, 2015). Under the additional assumption that the eigenvalues are distinct, we can also get the consistency and the asymptotic normality of the estimated eigenfunctions $\widehat{\boldsymbol{\phi}}_{r}(t)$, $r=1,\ldots,k$.
4 An illustration in sensory analysis
The dataset used for illustration deals with sensory experiments. It is well documented in Béno et al., (2023) and can be obtained from a public source. Computations were performed with the library MFPCA (see Happ,, 2022) in the R language R Core Team, (2024). All codes are available on Github, https://github.com/Chemosens/ExternalCode/tree/main/MFPCAWithCategoricalTrajectories.